# Monitoring Latent World States in Language Models with Propositional Probes

### v1

@@ Abstract
Language models are susceptible to bias, sycophancy, backdoors, and other tendencies that lead to unfaithful responses to the input context. Interpreting internal states of language models could help monitor and correct unfaithful behavior. We hypothesize that language models represent their input contexts in a latent world model, and seek to extract this latent world state from the activations. We do so with propositional probes, which compositionally probe tokens for lexical information and bind them into logical propositions representing the world state. For example, given the input context “Greg is a nurse. Laura is a physicist.”, we decode the propositions $\mathsf{WorksAs(Greg,nurse)}$ and $\mathsf{WorksAs(Laura,physicist)}$ from the model’s activations. Key to this is identifying a binding subspace in which bound tokens have high similarity (Greg $\leftrightarrow$ nurse) but unbound ones do not (Greg ↮$\not\leftrightarrow$ physicist). We validate propositional probes in a closed-world setting with finitely many predicates and properties. Despite being trained on simple templated contexts, propositional probes generalize to contexts rewritten as short stories and translated to Spanish. Moreover, we find that in three settings where language models respond unfaithfully to the input context—prompt injections, backdoor attacks, and gender bias— the decoded propositions remain faithful. This suggests that language models often encode a faithful world model but decode it unfaithfully, which motivates the search for better interpretability tools for monitoring LMs.
@@ 1 Introduction
Language models (LMs) appear to robustly understand their input contexts and flexibly respond to questions about them. Yet, they are susceptible to tendencies that lead to responses unfaithful to the input contexts: they are misled by irrelevant examples  [1, 2], subject to unintended tendencies  [3] and biases  [4, 5], and vulnerable to adversarial attacks on their prompts  [6] and training data  [7, 8].
Being able to interpret internal states of language models could help diagnose and correct language models when they become unfaithful [9, 10]. In particular, language models might internally represent the input context in a latent world model [11, 12]; when the LM is influenced by unfaithful tendencies, the latent world model could remain truthful even if the LM outputs falsehoods (Fig. 1). Thus, the state of the latent world model could provide more faithful information than the model outputs.
We extract the latent world state by introducing propositional probes, which probe the internal activations of the input context for logical propositions representing a world state. For example, the input context “The CEO and the nurse live in Thailand and France respectively.” would be represented by the propositions $\{\mathsf{LivesIn(CEO,France)},\mathsf{LivesIn(Nurse,France)}\}$ (Fig. 1).
@@
The key difficulty in predicting propositions is binding entities in predicates—the space of propositions is exponentially large, so we need to exploit compositionality. Our propositional probes probe individual tokens for lexical information, and bind them together to form propositions (Fig. 2). To determine binding we identify with a novel Hessian-based algorithm the “binding subspace” [13], a subspace of activation space in which bound tokens have similar activations. We find that the identified binding subspace causally mediates [14, 15, 16] binding. In Sec. 4 we describe and evaluate the Hessian algorithm, and in Sec. 5 we discuss how propositional probes are constructed.
We validate propositional probes in a closed-world setting about people and their attributes with a finite number of predicates and properties. We generate synthetic training data by populating templates with random propositions. To evaluate their generalizability, we rewrite the templated data into short stories using GPT-3.5-turbo, and translate the short stories into Spanish. We find that propositional probes generalize to these complex settings (Sec. 5.3).
We then apply propositional probes to three situations where language models behave unfaithfully, and show that they reflect the input context more faithfully than the outputs of the language models (Sec. 6). Specifically, we study cases where the language model is influenced by prompt injections, backdoors, and gender bias. Our results indicate that propositional probes, if scaled sufficiently, could be applied to monitoring language models at inference time.
Figure 1: Left: LMs construct representations of their inputs. They condition on these internal activations to answer queries (Prompting), but we extract symbolic propositions from them (Propositional Probes). Right: LMs may exhibit tendencies towards unfaithfulness, which we hypothesize will influence final responses to prompts but not the early-stage activations detected by our probes.
Figure 2: Left: Name probes (blue) and country probes (green) classify activations into either a name/country or a null value. Right: Activations have a content component and a binding component, such that bound activations have similar binding components. We use this to compose across tokens.
@@ 2 Related Work
Probing   Probes have been shown to be effective at decoding lexical information from language model activations in many domains  [17], such as color  [18], gender  [19], and space  [20]. These insights inform how we construct constituent domain probes. In addition, propositional probes have also been studied in autoregressive models trained on Othello games  [12] and in language models fine-tuned on closed-world Alchemy and TextWorld environments [11]. Lastly, probes have been used to study representations of factuality [21, 22] with respect to static knowledge learned from training; in contrast we study faithfulness with respect to dynamic knowledge described in context.
Monitoring Language Models   Researchers have argued that probes that extract latent knowledge in machine learning systems can provide monitoring tools necessary for usability and safety  [9, 10]. To benchmark monitoring methods, researchers have proposed an anomaly detection framework [23, 24]: probes should be trained only on simple trusted data and tested on hard, untrusted data. Our evaluations are consistent with this principle.
@@
Representations of binding   Researchers have long studied binding in connectionist models and human minds [25, 26, 27, 28]. In recent statistical language models, researchers have shown that representations of semantic roles and coreferences emerge from pretraining  [29, 30, 31]. We build on the “binding vectors” discovered in language model activations  [13, 32], which we describe later.
3 Preliminaries
Closed-world datasets   We consider a closed world about people and their attributes. Each person has a name, country of origin, occupation, and a food they like. We describe an instance of the world with a set of propositions, such as $\mathsf{LivesIn(Carol,Italy)}$. Each proposition consists of one of three predicates (“LivesIn”, “WorksAs”, “LikesToEat”), each of which binds a name to a country, occupation, or food respectively.
In greater detail, we construct four domains, denoted as $\mathcal{D}_{0},\dots,\mathcal{D}_{3}$, which correspond to names, countries, occupations, and foods respectively. Each domain $\mathcal{D}_{k}$ is a set of $|\mathcal{D}_{k}|$ values, e.g. $\mathcal{D}_{0}$ is a set of names. Then, a proposition consists of a predicate (either “LivesIn”, “WorksAs”, or “LikesToEat”), a name $E\in\mathcal{D}_{0}$, and an attribute $A$ from $\mathcal{D}_{1},\mathcal{D}_{2},$ or $\mathcal{D}_{3}$ respectively.
In our setting, the predicates correspond one-to-one to the attribute domains, which implies that a proposition can be determined by a value from the names domain and a value from one of the three attribute domains—the predicate can be inferred from which domain the attribute came from. Our method utilizes this observation. In addition, for brevity we sometimes drop the predicate, e.g. $\mathsf{(Carol,Italy)}$ instead of $\mathsf{LivesIn(Carol,Italy)}$.
To create the input context, we generate random propositions about 2 people. For each person, we generate a random name, country, occupation and food, and instantiate the three propositions about them. We make sure that the two people do not share any attribute. These propositions are formatted in a template to produce the synth dataset, rewritten using GPT-3.5-turbo into a short story (para), and then translated into Spanish (trans). See Fig. 3 for an example and Appendix A for details.
Figure 3: Dataset creation: Sets of propositions about 2 people are generated randomly. Each set is formatted by a template (synth), rewritten into a story (para), and translated into Spanish (trans).
Domain Probes   For every domain, we train a domain probe that linearly classifies activations at individual token positions into either a value in the domain, or a null value, indicating that none of the values is represented. For example, Fig. 2 (left) shows the outputs of the name probe and the country probe. In greater detail, the language model converts a context with $S$ tokens, $T_{0},\dots T_{S-1}$ into internal activations $Z_{0},\dots,Z_{S-1}$. For every domain $\mathcal{D}_{k}$, we train a probe $P_{k}$ that takes as input an activation $Z_{s}$ at any token position $s$, and outputs a value in $\mathcal{D}_{k}\cup\{\bot\}$. In Sec. 5.1 we describe how we use simple linear probes to implement domain probes.
Binding subspace   We use the binding subspace to compose values identified by domain probes to form propositions. Prior work  [13] suggested that activations can be linearly decomposed into two components: a content vector and a binding vector (Fig. 2, right); moreover, two activations are bound if and only if their binding vectors are similar. Additionally, they suggested that binding vectors may form a binding subspace, which we identify in this work using a Hessian-based algorithm (Sec. 4.1). We use the binding subspace to form a binding similarity metric between token activations that informs how domain probe results are bound together in propositions.
Models   Throughout the paper we use the Tulu-2 13 billion parameter model [33], an instruction-tuned version of the Llama 2 base model [34]. Our evaluations on Llama-2-13b-chat did not differ significantly. The models have $n_{\text{layers}}=40$ layers and $d_{\text{model}}=5120$ embedding dimensions.
4 Identifying the binding subspace
In this section, we present a Hessian-based algorithm for identifying the binding subspace (Sec. 4.1), which is loosely inspired by the use of the Jacobian in an unrelated task [35]. To evaluate the binding subspace, we perform quantitative causal interventions as well as qualitative analysis (Sec. 4.2).
4.1 Hessian-based algorithm
In this section, we first motivate the Hessian-based algorithm with a thought experiment. We then describe the algorithm, which produces a matrix $H$ that characterizes the binding subspace. Finally, we describe how the binding subspace is extracted from $H$, and how we use $H$ to construct a binding similarity metric $d(Z_{x},Z_{y})$ that measures how bound any two activations $Z_{x},Z_{y}$ are.
Let us imagine there really is a binding subspace where two activations $Z_{x}$ and $Z_{y}$ are bound only if $Z_{x}^{\top}HZ_{y}$ is large for some low rank $H$. Then, suppose we take two initially unbound activations $Z_{x}$ and $Z_{y}$ and perturb the first in direction $x$ and the latter in direction $y$. The binding strength should increase only if directions $x$ and $y$ align under $H$. In fact, if there is a measure of binding strength $F(x,y)$ that is bilinear in $x,y$, then $H$ is precisely the second-derivative $\nabla_{x}\nabla_{y}F(x,y)$.
To instantiate this, we need a way of creating unbound activations, and also of measuring binding strength. To create unbound activations, we construct activations that are initially bound and subtract away the binding information so that the binding vectors become indistinguishable. Specifically, we consider a context with two propositions $\{(E_{0},A_{0}),(E_{1},A_{1})\}$, and let $Z_{E_{i}}$ and $Z_{A_{i}}$ for $i=0,1$ be the activations for $E_{i}$ and $A_{i}$ respectively. Prior work [13] showed we can decompose the activation $Z_{E_{0}}$ into a content part $f_{E_{0}}$ and a binding part $b_{E_{0}}$, and likewise for $E_{1},A_{0},$ and $A_{1}$. We follow the averaging procedure in prior work (explained later) to estimate differences between the binding vectors, $\smash{\Delta_{E}\triangleq b_{E_{1}}-b_{E_{0}},\Delta_{A}\triangleq b_{A_{1}}-b_{A_{0}}}$. Then, to erase binding information, we add $0.5\Delta_{E}$ and $0.5\Delta_{A}$ to $Z_{E_{0}}$ and $Z_{A_{0}}$, and subtract the same quantities from $Z_{E_{1}}$ and $Z_{A_{1}}$. This moves the binding vectors to their midpoint and so should cause them to be indistinguishable from each other.
To measure binding strength, we append to the context a query that asks for the attribute bound to either $E_{0}$ or $E_{1}$. For example, if the attributes are the countries of residence of the entities, we would ask “Which country does $E_{0}$/$E_{1}$ live in?”. We measure the probability assigned to the correct answer ($A_{0}$ and $A_{1}$ respectively), and take the average over the two queries as the measure of binding strength $F(x,y)$, where $x$ and $y$ are the perturbations added to the unbound entity and attribute activations. Therefore, when all activations are unbound ($x=y=0$), the model takes a random guess between the two attributes, and so $F(0,0)=0.5$. If $x$ and $y$ are aligned under $H$, we expect $F(x,y)>0.5$.
The matrix $H$ is obtained as the second-derivative $\nabla_{x}\nabla_{y}F(x,y)$. For tractability, we parameterize $x$ and $y$ so that they are shared across layers, i.e. $\smash{x,y\in\mathbb{R}^{d_{\text{model}}}}$ instead of $\smash{\mathbb{R}^{d_{\text{model}}\times n_{\text{layers}}}}$.
To obtain the low-rank binding subspace from $H$, we take its singular value decomposition $H=USV^{\top}$, where $S$ is a diagonal matrix with decreasing entries. We expect the binding subspace to be the top $k$-dimensional subspaces of $U$ and $V$, $U_{(k)},V_{(k)}\in\mathbb{R}^{d_{\text{model}}\times k}$, for a relatively small value of $k$.
Binding similarity metric   To turn the binding subspace $U_{(k)}$ into a similarity metric between two activations $Z_{s}$ and $Z_{t}$, we take the activations $Z_{s}^{(l)},Z_{t}^{(l)}$ at a certain layer $l$, project them into $U_{(k)}$, and compute their inner product under the metric induced by $S$:
$d(Z_{s},Z_{t})\triangleq Z^{(l)\top}_{s}U_{(k)}S_{(k)}^{2}U_{(k)}^{\top}Z^{(l)}_{t}.$ (1)
Estimating $\Delta_{E}$ and $\Delta_{A}$   Following prior work [13], to estimate $\Delta_{E}=b_{E_{1}}-b_{E_{0}}$ we take the average difference $Z_{E_{1}}-Z_{E_{0}}$ across 200 contexts with different values of $E_{0}$ and $E_{1}$ with the hope that the content dependent vectors $f_{E_{0}},f_{E_{1}}$ cancel out (likewise for $\Delta_{A}$).
4.2 Evaluations of the Hessian-based algorithm
In this section, we first show quantitatively that the Hessian-based algorithm provides a subspace that causally mediates binding, and that this subspace generalizes to contexts with three entities even though the Hessian was taken for two-entity contexts. We then qualitatively evaluate our binding subspace by plotting the binding similarity (Eq. 1) for a few input contexts.
@@ 4.2 Evaluations of the Hessian-based algorithm
$Z_{E_{0}}^{(l)}$ $\leftarrow(I-P)Z_{E_{0}}^{(l)}+PZ_{E_{1}}$ (2)
$Z_{E_{1}}^{(l)}$ $\leftarrow(I-P)Z_{E_{1}}^{(l)}+PZ_{E_{0}}^{(l)},$ (3)
with $P=U_{(k)}U_{(k)}^{\top}$. If $U_{(k)}$ correctly captures binding information, then we expect the binding information to have swapped for $E_{0}$ and $E_{1}$.
We apply the interchange intervention with different dimensions $k$ for entities using $U_{(k)}$, and for attributes using subspace $V_{(k)}$ with analogous interventions (Eq. (2), (3)) on $Z_{A_{i}}$ and $Z_{A_{j}}$ instead (Fig. 4). As a random baseline, we take the SVD of a random matrix instead of the Hessian. As an additional baseline, we implement implement Distributed Alignment Search (DAS) [37]. Specifically, we use gradient descent to find a fixed dimensional subspace that enables interchange interventions between the two entities in two-entity contexts, for various choices of subspace dimension.
As a skyline, we evaluate the two-dimensional subspace spanned by the differences in binding vectors of the three entities. We obtain these difference vectors similarly as $\Delta_{E}$ and $\Delta_{A}$ used in the computation of the Hessian: we take samples of $Z_{E_{0}},Z_{E_{1}},$ and $Z_{E_{2}}$ across 200 contexts with different values of entities, average across the samples, and take differences between the three resultant mean activations. We consider this a skyline because this subspace is obtained from three entities’ binding vectors, whereas the Hessian and DAS are computed using contexts with only two entities.
The results show that the top 50 dimensions of the Hessian (out of $5120$) are enough to capture binding information. Moreover, despite being obtained from contexts with two entities, the subspace correctly identifies the direction of the third binding vector, in that it enables the swaps “0-2” and “1-2”. In contrast, random subspaces of all dimensions fail to switch binding information without also switching content. Further, while DAS finds subspaces that successfully swaps the first two binding vectors, they do not enable swaps between the second and third binding vectors (1-2). Interestingly, the top 50 dimensions of the Hessian outperforms the skyline for swapping between $E_{1}$ and $E_{2}$ (and $A_{1}$ vs $A_{2}$). This could be due to the fact that the skyline subspace is obtained from differences in binding vectors, which could contain systematic errors that do not contribute to binding, or perhaps due to variability in binding vectors in individual inputs. We discuss details in Appendix D.
Figure 4: The accuracy of swapping binding information in name (attribute) activations by projecting into $U_{(k)}$ ($V_{(k)}$) against $k$ in a context with 3 names and 3 attributes. We test the subspaces from the Hessian (blue), a random baseline (orange), and a skyline subspace obtained by estimating the subspace spanned by the first 3 binding vectors. We perform all 3 pairwise switches: 0-1 represents swapping the binding information of $E_{0}$ and $E_{1}$ ($A_{0}$ and $A_{1}$), and so on.
Qualitative evaluations   We visualize the binding metric in various contexts by plotting pairwise binding similarities (Eq. 1) between token activations. Specifically, on a set of activations $Z_{0},\dots,Z_{S-1}$, we compute the matrix $M_{st}=d(Z_{s},Z_{t})$. We show the similarity matrix for three contexts in Fig. 5. We first evaluate an input context in which the entities are bound serially: the first sentence binds $E_{0}$ to $A_{0}$, and the second binds $E_{1}$ to $A_{1}$. To ensure that the binding subspace is picking up on binding, and not something spurious such as sentence boundaries, we evaluate an input context in which entities are bound in parallel: there is now only one sentence that binds $E_{0}$ and $E_{1}$ to $A_{0}$ and $A_{1}$ respectively. In both the serial context (Fig. 5 left) and parallel context (Fig. 5 middle), the activations are clustered based on which entity they refer to. In addition, we plot the similarity matrix for a context with three entities (Fig. 5 right). Interestingly, the binding metric does not clearly discriminate between the second and third entities even though the interchange interventions showed that the binding subspace captures the difference in binding between them. This suggests that the 50-dimensional binding subspace obtained from the Hessian may either contain spurious non-binding directions or only incompletely captures the binding subspace. Thus, our current methods may be too noisy for contexts with more than two entities. In Appendix E, we use similar plots to show that coreferred entities share the same binding vectors.

### v2

@@ Abstract
Language models (LMs) are susceptible to bias, sycophancy, backdoors, and other tendencies that lead to unfaithful responses to the input context. Interpreting internal states of LMs could help monitor and correct unfaithful behavior. We hypothesize that LMs faithfully represent their input contexts in a latent world model, and we seek to extract these latent world states as logical propositions. For example, given the input context “Greg is a nurse. Laura is a physicist.”, we aim to decode the propositions $\mathsf{WorksAs(Greg,nurse)}$ and $\mathsf{WorksAs(Laura,physicist)}$ from the model’s internal activations. To do so we introduce propositional probes, which compositionally extract lexical concepts from token activations and bind them into propositions. Key to this is identifying a binding subspace in which bound tokens have high similarity ($\mathsf{Greg}\leftrightarrow\mathsf{nurse}$) but unbound ones do not ($\mathsf{Greg}\not\leftrightarrow\mathsf{physicist}$). Despite only being trained on linguistically simple English templates, we find that propositional probes generalize to inputs written as short stories and translated to Spanish. Moreover, in three settings where LMs respond unfaithfully to the input context—prompt injections, backdoor attacks, and gender bias— the decoded propositions remain faithful. This suggests that LMs often encode a faithful world model but decode it unfaithfully, which motivates the search for better interpretability tools for monitoring LMs.
@@ 1 Introduction
Language models (LMs) may produce responses unfaithful to the input context in many situations: they can be misled by irrelevant examples  (Anil et al., 2024; Halawi et al., 2023), learn unintended tendencies  (Sharma et al., 2023) and biases  (Blodgett et al., 2020; Liang et al., 2022), and are vulnerable to adversarial attacks on their prompts  (Perez & Ribeiro, 2022) and training data  (Wallace et al., 2020; Hubinger et al., 2024). This makes it important to be able to identify and correct when language models are unfaithful.
To address unfaithfulness, one promising approach is to interpret LMs’ internal states (Christiano et al., 2021; Viégas & Wattenberg, 2023). A common hypothesis is that LMs internally represent the input context in a latent world model (Li et al., 2021; 2022), which could remain faithful even if the LM outputs falsehoods (Mallen & Belrose, 2023). This might happen, for instance, if unfaithful tendencies affect how the model behaves in accordance to its beliefs, but not the beliefs themselves.
In this work, we introduce propositional probes, which extract symbolic latent world states as logical propositions; we then show that these latent world states remain faithful in three adversarial settings. As an example of propositional probes, consider the input context “Alice lives in Laos. Bob lives in Peru.” We would extract from internal activations the propositions $\{\mathsf{LivesIn(Alice,Laos)}$, $\mathsf{LivesIn(Bob,Peru)}\}$ In three adversarial settings, namely prompt injections, backdoors, and gender bias, we find that these extracted latent world states are more faithful than model outputs.
@@
Our probes work by extracting lexical concepts from token activations, then binding them together to form propositions (Fig. 1) (Sec. 4). Since the space of propositions is exponentially large, our probes exploit compositionality by building propositions from their lexical constituents (e.g. composing $\mathsf{LivesIn(Alice,Laos)}$ from $\mathsf{Alice}$ and $\mathsf{Laos}$)). Key to enabling this is the “binding subspace” (Feng & Steinhardt, 2023), a subspace of activation space in which bound tokens (such as Alice and Laos above) have similar activations. To identify this subspace, we develop a novel Hessian-based algorithm (Sec. 5) and demonstrate that the identified binding subspace causally mediates binding (Geiger et al., 2021a; Vig et al., 2020; Pearl, 2022) (Sec. 5.3).
Our propositional probes generalize well to complex and adversarial settings, despite being trained only on linguistically simple data. We create the training data by filling simple templates with random ground-truth propositions drawn from a finite universe of predicates and objects, and the test data by rewriting the train data into short stories and translating them into Spanish with GPT-3.5-turbo (Sec. 3). We find that propositional probes generalize to these complex settings, achieving a Jaccard Index of within 10% of a prompting skyline (Sec. 6.1). Further, in three adversarial settings (prompt injections, backdoors, and gender bias), our probes produce more faithful propositions than what the model outputs (Sec. 6.2).
Overall, our results indicate that identifying compositional structures in LM activations is a promising approach towards building monitoring systems for LMs. This is particularly pertinent as LMs are increasingly deployed as autonomous agents. To this end, we hope that our work could stimulate further research into extracting world models of greater complexity, such as by modeling role-filler binding (Smolensky, 1990) and state changes (Kim & Schuster, 2023).
[…]
@@ 2 Related Work
Probing   Our propositional probes compose smaller lexical probes together; the constituent probes have been well studied in many domains (Mikolov et al., 2013), such as color  (Abdou et al., 2021), gender  (Bolukbasi et al., 2016), and space  (Gurnee & Tegmark, 2023). In addition, probing for propositional beliefs has been studied for the Othello (Li et al., 2022), Alchemy, and TextWorld environments (Li et al., 2021). Our approach differs by exploiting compositionality, which enables our method to leverage the large body of existing probing literature.
@@
Representations of binding   Researchers have long studied binding in connectionist models and human minds (von der Malsburg, 1981; Feldman, 1982; 2013; Treisman, 1996). In recent LMs, researchers have shown that representations of semantic roles and coreferences emerge from pretraining  (Tenney et al., 2019; Belinkov et al., 2020; Peters et al., 2018). We build on the “binding vectors” discovered in language model activations  (Feng & Steinhardt, 2023; Prakash et al., 2024).
3 Task definition and preliminaries
In this section, we formally define the task that propositional probes solve, and discuss how we evaluate propositional probes.
@@ 3 Task definition and preliminaries
Easy-to-hard evaluations   To evaluate propositional probes we train our probes on simple inputs and test their generalization on hard, potentially adversarial, inputs. Prior works have used this easy-to-hard evaluation paradigm (Roger et al., 2023; Mallen & Belrose, 2023) to measure the usefulness of probes as monitoring methods. Specifically, we create three versions of datasets; the propositional probes are trained only using the simplest version (synth), and evaluated on the heldout versions (para, trans). Further, in Sec. 6.2 we introduce additional distribution shifts during evaluation that adversarially changes the model’s output behavior, and test if the probes can remain faithful.
Datasets   We create three datasets of natural language passages of increasing complexity, namely synth, para, and trans, each labelled with ground-truth propositions. To do so, we first generate random propositions, which are then formatted in a template to produce the synth dataset, and augmented using GPT-3.5-turbo to produce diverse and complex para and trans datasets (Fig. 2).
In more detail, each input instance describes two people, each of whom has a name, country of origin, occupation, and a food they like. We model this information as a set of propositions such as $\mathsf{LivesIn(Carol,Italy)}$. Formally, our closed world consists of four domains $\mathcal{D}_{0},\dots,\mathcal{D}_{3}$, which are sets of names, countries, occupations, and foods respectively. There are three predicates, $\mathsf{LivesIn}$, $\mathsf{WorksAs}$, and $\mathsf{LikesToEat}$, each of which binds a name $E\in\mathcal{D}_{0}$ to an attribute from one of the three attribute domains $\mathcal{D}_{1},\mathcal{D}_{2},$ or $\mathcal{D}_{3}$ respectively. Note that because of the one-to-one correspondence between predicates and attribute domains, the predicate of a proposition can be inferred by which domain the attribute is from. Our method utilizes this observation. In addition, for brevity we sometimes drop the predicate, e.g. $\mathsf{(Carol,Italy)}$ instead of $\mathsf{LivesIn(Carol,Italy)}$.
To create the input context, we generate six random propositions about two people by sampling without repetition two values from each domain. These propositions are formatted in a template to produce the synth dataset, rewritten using GPT-3.5-turbo into a short story (para), and then translated into Spanish (trans). See Fig. 2 for an example and Appendix A for details.
Figure 2: To create our datasets, we first generate sets of random propositions about two people. Each set is formatted with a template (synth), rewritten into a story (para), and translated into Spanish (trans). We train probes to predict propositions from the easy synth dataset, and test probes on the hard para and trans datasets.
Models   We use the Tulu-2-13b model (Ivison et al., 2023), an instruction-tuned version of Llama 2 (Touvron et al., 2023) with $n_{\text{layers}}=40$ layers and $d_{\text{model}}=5120$ embedding dimensions.